### v1

# Laser mode-hopping assisted all-optical single beam pulsed atomic magnetometer

Abstract
We demonstrate an all-optical single beam pulsed atomic magnetometer assisted by laser mode-hopping in a distributed Bragg reflector (DBR) laser. We implement a temporal sequence of the laser current; sinusoidal current modulation including the laser mode-hop current for synchronous optical pumping and the following constant current for paramagnetic Faraday rotation measurements to probe the free induction decay (FID) of transverse 87Rb spin polarization. Repetitive sudden frequency shifts of 20 GHz around the pressure-broadened 87Rb spectra originating from laser mode-hopping enables discontinuous optical pumping modulation with a large depth which enhances transverse spin polarization. We achieve a sensitivity of 3.77 pT/Hz1/2 in a magnetic field of 14 $\mu$T, limited by the performance of the frequency counter. The Cramer-Rao lower bound (CRLB) of the sensitivity due to the non-magnetic noise such as photon shot-noise is 191 fT/Hz1/2. Our approach based on laser mode-hopping can be applied to miniaturization of all-optical atomic magnetometers with sub-pT/Hz1/2 sensitivities.
I Introduction
Laser mode-hopping refers to a jump of the lasing mode, which stems from the relative frequency shift of longitudinal cavity modes to the gain curve originating from temperature or injection current changes. It is generally considered to be problematic in laser spectroscopy and telecommunication because it accompanies unwanted intensity fluctuations and large sudden frequency shifts. Many researches have been performed to extend the mode-hop free range to facilitate continuous wavelength tuning gong2014mode ; dutta2012mode ; fuhrer2008extension . In the aspect of understading underlying physics of mode-hopping, there also have been attempts to study dynamics of laser mode-hopping kundu2018ultrafast ; butler2017direct ; happach2020effect .
A large sudden frequency shift feature during laser mode-hopping, however, may be useful in some applications; in Bell-Bloom type atomic magnetometer experiments where the modulation of optical pumping rate to induce spin-precession is involved, large sudden frequency shifts during mode-hopping offers enhancement of spin-polarization by increasing the modulation depth of optical pumping. In order to build up large transverse spin polarization in all-optical Bell-Bloom type atomic magnetometers, a variety of techniques including sinusoidal modulation of the laser frequency has been explored kimball2017situ ; jimenez2009sensitivity . Some advancements in nonlinear magneto-optical rotation (NMOR) magnetometer were presented based on pulse-shaped amplitude modulation of beam, taking advantage of the discontinuous modulation gawlik2006nonlinear ; gerginov2017pulsed . Obtaining a discontinuous lineshape in the modulation function, however, requires cumbersome optics like acousto-optic modulators (AOM) or intricate techniques to generate square-wave RF magnetic field ranjbaran2017effects ; hunter2018free .
Frequency switching by means of mode-hopping, on the other hand, only requires weak sinusoidal modulation of the laser diode (LD) injection current to create a large discontinuous optical pumping modulation. Such an effectiveness can be estimated from the optical pumping rate formula. The optical pumping rate is a function of laser intensity and detuning, given by $R\propto I/(\Delta\nu^{2}+(\Gamma_{L}/2)^{2})$ where $I$ is the laser intensity, $\Delta\nu$ is the laser detuning, and $\Gamma_{L}$ is the line broadening. For example, when the frequency shift during a mode-hopping is $2\Gamma_{L}$, the optical pumping rate of $\Delta\nu=2\Gamma_{L}$ is 17 times smaller than that of $\Delta\nu=0$. To acheive the same modulation depth by using an intensity modulation, the laser intensity have to be changed from $I$ to $I/17$. To obtain such a large modulation depth an external light modulator, for example, AOM or electro-optic modulator (EOM), can be applied. For mininarization of the device, direct injection current modulation into the LD can be adopted instead of using an external light modulator. Considering the linear frequency shift of the LD as a function of injection current, in our case, -1 GHz/mA in the mode-hop free range, an amplitude of current modulation of 20 mA is required to reach a 20 GHz detuning corresponding to a double of $\Gamma_{L}=10$ GHz, a typical line broadening of the Rb D1 line in nitrogen buffer gas. Such a detuning can be easily obtained by only a few mA when laser mode-hopping is adopted. Moreover, the injection current into the LD is not needed to be discontinuous. A continuous laser modulation with a small modulation depth leads to a large depth of discontinuous optical pumping modulation when laser mode-hopping is applied. It is an advantage of laser mode-hopping assisted optical pumping because a large amplitude of discontinuous laser injection current can be a source of the unwanted magnetic field noise.
Frequency switching feature can be also helpful for implementing single beam all-optical atomic magnetometers based on FID measurements. We can easily shift the laser frequency to be far detuned from or to be resonant on the atomic resonance by utilizing laser mode-hopping. A temporal sequence of the laser current containing the mode-hop current enables FID measurements. In the on-resonance period where the laser is resonant on the atomic transition, the atomic spin is polarized along the laser beam direction. In the following off-resonance period where the laser is far detuned from the atomic transition due to the mode-hopping, optical pumping ceases and the FID of the transversal atomic spin is induced. The far-detuned laser undergoes the paramagnetic Faraday rotation depending on the transversal atomic spin which precesses with the Larmor frequency.
Usually, single beam atomic magnetometers utilize a magnetic field modulation technique using a coil liu2020single ; ding2020single ; oelsner2022integrated . Their structure are very simple but they can suffer from magnetic field interference when forming a gradiometer or a magnetometer array. To diminish such the effect of magnetic interference all-optical atomic magnetometers have been developed widely. All-optical vector atomic magnetometer was reported patton2014all . An all-optical atomic magnetometer with a co-linear configuration to measure the biomedical magnetic field was implemented by placing the pump and probe lasers at a slight angle within the same window limes2020portable . Single beam all-optical atomic magnetometers based on an external light modulator were studied petrenko2022towards ; hunter2018free . Even, all-optical single beam atomic magnetometer based on direct current modulation to the vertical-cavity surface-emitting laser (VCSEL) was already demonstrated with a sensitivity level of 10 pT/Hz1/2 jimenez2009sensitivity ; hunter2018free . However, the low power of the VCSEL is insufficient to obtain a significant optical pumping rate $R(t)$ so that the sensitivity is limited. Our strategy for single beam all-optical atomic magnetometer without the use of an external light modulator is that adopting a DBR laser instead of a VCSEL to acheive higher optical pumping rate and applying current sequence including a laser mode-hopping, which consists of the optical pumping stage and the following probing stage to measure the FID of macroscopic 87Rb spin.
In this paper, we demonstrate an all-optical single beam pulsed atomic magnetometer utilizing a mode-hopping of a DBR laser. We employ mode-hopping to selectively switch the laser frequency between on-resonance and off-resonance with the 87Rb D1 line so that a sequence of optical pumping and the FID probing is realized. We observe the FID as a function of offset, modulation amplitude of current modulation and analyze the discontinuous optical pumping effect. Additionally, we measure the sensitivity of our magnetometer and compare it to the CRLB due to the non-magnetic noise which shows the reachable sensitivity of our magnetometer.
II Model for discontinuous optical pumping
Figure 1: Illustration of the concept of frequency modulation assisted by laser mode-hopping. A sinusoidal modulation of laser injection current is converted to a discontinuous output frequency profile. The red and blue area represent the adjacent cavity modes in a DBR laser. A small variation around the mode-hop current results in a huge shift on laser output frequency. The frequencies of these modes correspond to the resonant and off-resonant condition of pressure broadened 87Rb absorption profile, respectively. The repetitive jump between cavity modes induces a discontinuous synchronous optical pumping.
A discontinuous optical pumping profile is achievable by incorporating a simple sinusoidal modulation of the laser injection current with assistance of a mode-hopping (see Figure 1). When the sinusoidal current modulation contains the mode-hop current, the laser frequency jumps by the frequency difference between cavity modes. Assumming that the frequency of the mode 1 (or the mode 2) is resonant on (or far-detuned from) the atomic resonance, the optical pumping rate, $R\propto I/(\Delta\nu^{2}+(\Gamma_{L}/2)^{2})$, shows steep changes as shown in Fig. 1. The change on optical pumping rate can be regarded to be sufficiently sudden, in other words, discontinous because the swithcing time of mode-hopping given by an order of 10 ns is much shorter than the Larmor period of 2.85 $\mu$s in a magnetic field of $50~{}\mu$T butler2017direct .
In order to analyze the effect of the discontinuous optical pumping induced by a mode-hopping, we investigate the macroscopic spin Bloch equation including the optical pumping modulation term, $R(t)$. The Bloch equation model, widely used for atomic magnetometers, is suitable for modeling arbitrary modulations distorted by mode-hopping grujic2013atomic . The Bloch equation is given by
$\frac{d\langle\vec{S}\rangle}{dt}=\gamma\langle\vec{S}\rangle\times B_{y}\hat{y}-\Gamma_{re}\langle\vec{S}\rangle+R(t)(\frac{\vec{s}}{2}-\langle\vec{S}\rangle).$ (1)
Here, $\vec{s}=(e^{*}\times e)/i$ represents the photon spin vector where $e$ represents a unit polarization vector in the yz plane so that $\vec{s}=s\hat{x}$. $\Gamma_{re}$ is the spin relaxation rate mainly contributed by the spin exchange collision between 87Rb atoms. $R(t)$ is the optical pumping rate for unpolarized atoms, depending on the laser detuning from the $D_{1}$ resonance and the laser intensity vasilakis2011precision . $\gamma$ is the gyromagnetic ratio. We consider the magnetic field is directed in the $\hat{y}$ axis.
[…]
$S_{0}\approx\frac{(R_{1}+R_{-1})|s|}{R_{0}}\frac{1}{4(1+\Gamma_{re}/R_{0})}.$ (2)
In the case of sinusoidal optical pumping modulation described by $R(t)=A+B\cos(\omega t)$, the value of $(R_{1}+R_{-1})/R_{0}$ is calculated as $B/A$, which can not exceed 1 because $R(t)$ is always positive, i.e., $A\geq B$. The maximum $S_{0}$ is 1/4 in the sinusoidal optical pumping case when $R_{0}=R_{1}$ and $R_{0}\gg\Gamma_{re}$. A discontinuous $R(t)$ (e.g. square-wave function), on the other hand, allows for $(R_{1}+R_{-1})/R_{0}>1$.
In the case of square wave optical pumping modulation, $R_{0}$ varies accordingly to the duty cycle $d$ when the maximum value of $R(t)$ is fixed. Assuming the maximum optical pumping rate of the square wave optical pumping modulation is $R_{m}$, then $S_{0}$ can be expressed by
$S_{0}\approx\frac{|s|\text{sinc}(\pi d)}{2(1+\frac{\Gamma_{re}}{R_{m}d})}.$ (3)
The maximum $S_{0}$ is 1/2 when $d\rightarrow 0$ in the regime that $\Gamma_{re}\ll R_{m}d$. This implies that a delta-function like optical pumping induces the largest $S_{0}$ as long as the peak intensity of a laser pulse is strong enough to satisfy $R_{m}\gg\Gamma_{re}/d$. The reason why the delta-function like optical pumping is optimal can be described in terms of $\vec{s}$ and $\langle\vec{S}\rangle$. In the Bell-Bloom configuration, optical pumping not only creates the Rb spin polarization along the direction of the pumping beam but also relaxes the transverse spin. When the direction of $\vec{s}$ does not coincide with $\langle\vec{S}\rangle$, optical puming destructs the transverse spin precession and aligns the atomic spin with the direction of the pumping beam. It ceases to serve as a transverse spin relaxation mechanism only if $\langle\vec{S}\rangle$ is oriented in the same direction of $\vec{s}$. As a result, we can conclude that the delta-function like optical pumping is optimal. In the regime that the optical pumping rate $R_{m}$ is not much larger than $\Gamma_{re}/d$, the nonzero $d$ can be optimal. For example, the weak pumping condition that $\Gamma_{re}=2\pi\times 685$ Hz, $R_{m}=2\pi\times 300$ Hz, and $|s|=1/2$, the optimal duty cycle is caculated to be around 50 %. In this case, $S_{0}$ of the discontinuous optical pumping of a 50 % duty cycle is $4/\pi$ times larger than that of the sinusoidal optical pumping modulation if the maximum optical pumping rates are same.
III Experimental Setup
Figure 2: Experimental setup for the single beam all-optical magnetometer. DBR laser: Distributed Bragg Reflector laser, PBS: Polarization Beam Splitter, QWP: Quarter Wave Plate, HWP: Half Wave Plate, BPD: Balanced Photodetector. A burst sine wave modulation to the DBR laser relizes pulsed operation in 1 kHz. Red and blue regions show pumping and probing stage of pulsed magnetometer, respectively.
Figure 3: (a) 87Rb D1 absorption spectrum in the atomic vapor cell. The grey dots indicate the absorption of the linearly polarized laser beam. The black solid line depicts the fitting curve. The Full Width at Half Maximum linewidth $\Gamma_{L}$ is 10.5 GHz which implies that the pressure of $\textrm{N}_{2}$ in our cell is about 580 Torr. The red and blue solid lines depict the absorption spectrum of hyperfine transitions from $F_{g}=1$ and $F_{g}=2$, respectively. The black arrows show how the laser frequency detuning changes as the laser current grows. (b) Laser detuning as a function of LD injection currents. The mode-hop current is 67.1 mA. (c) Relative Faraday rotation angle (black solid line) and transverse spin relaxation rate (red solid line) $\Gamma_{re}$ as a function of LD injection current. The injection current below the mode-hop current induces a larger Faraday rotation angle and transverse spin relaxation rate because the corresponding laser frequency is within the 87Rb D1 absorption spectrum. (d) Relative optical power of the laser beam after passing through the atomic vapor cell as a function of LD injection current. The laser injection current above the mode-hop current gives higher beam power because the off-resonant light expriences weak absorption.
The experimental setup is shown in Figure 2. A DBR laser (Photodigm 794.978 nm series with TOSA package) is used as a light source. Mode-hopping in the DBR laser occurs under fixed condition due to it’s monolithic nature, allowing the jump to the output frequency controllable. Two main factors influence the frequency tuning of the DBR laser output: the temperature of the LD heat sink and the laser injection current. When modulating injection current with a stabilized laser temperature, a slight alteration in laser current triggers a transition between two adjacent cavity modes as illustrated in Fig. 1. The transition of lasing mode near the mode-hop current results in a significant jump in the laser output frequency, which was almost 20 GHz in our case. A sinusoidal modulation of LD injection current with aid of a mode-hopping, therefore, is converted to a discontinuous laser output frequency profile without any extra optics.
The linearly polarized laser beam is first transformed into an elliptically polarized beam after passing through a QWP. The optimal QWP angle to increase the signal is 25 and the correspoding $|s|$ is about 0.67 (see the appendix). The laser beam serves both as the probing and pumping beam. It illuminates a $5\times 5\times 5$ m3 cubic cell which contains 87Rb vapor and $\mathrm{N_{2}}$ gas at a pressure of 580 Torr (see Figure 3 (a)).
The home-made atomic gas cell is heated up to 103 C using a dual-layer structured pair of AC heaters designed to negate magnetic field noise yim2018note . After passing through the atomic gas cell, which is inside a two layered $\mu$-metal shield, the laser beam goes through a HWP. A polarimeter, consisting of a PBS and a BPD, measures the degree of polarization rotation of elliptically polarized beam induced by the paramagnetic Faraday rotation effect. The rotation angle of the polarization is proporitonal to $\Delta\nu/[\Delta\nu^{2}+(\Gamma_{L}/2)^{2}]$ where $\Delta\nu$ is the laser detuning in the probing stage.
Our pulsed magnetometer alternates pumping and probing mode at a repetition rate of 1 kHz. The laser injection current is modulated by a burst sine wave generated by a function generator (Siglent’s SDG2082X), realizing a pulsed operation of magnetometer. In the pumping stage, depicted as red region in Fig. 2, the LD injection current is modulated at the Larmor frequency of ${}^{87}Rb$ (which is 100 kHz in our case), incorporating a mode-hop current into the modulation of the LD injection current. The mode-hop current is 67.1 mA when the TEC temperature of the LD is set to 32.05 C. When the LD injection current is above the mode-hop current, the laser output frequency deviates from the Rb resonance line as laser detuning becomes almost 20 GHz, shown as an arrow in Fig. 3. On the other hand, when the LD current is below the mode-hop current, the laser output frequency gets close to the $F_{g}=1$ transition line as laser detuning becomes almost 0 GHz. Due to the pressure broadening of 10.5 GHz, the laser beam also weakly induces the transitions from $F_{g}=2$.
In the probing stage, illustrated by the blue region in Fig. 2, the sinusoidal modulation of the LD injection current stops and injection current is fixed to a certain value at which the laser frequency is off-resonant. The detuned laser beam undergoes the paramagnetic Faraday rotation effect where the polarization rotation angle is proportional to the atomic spin projected in the direction of the laser beam propagation. The laser beam with rotated polarization is directed into a polarimeter. A damped sine signal on which the FID is imprinted can be obtained from the BPD. It can be described by
$V(t,\Delta\nu)=V_{0}(\Delta\nu)e^{-\Gamma_{re}t}\cos(\omega_{0}t+\phi)$ (4)
where $V_{0}(\Delta\nu)$ is the FID amplitude which is proportional to $S_{0}$ multiplied by beam intensity after absorption $I_{0}$ and depends on the laser detuning $\Delta\nu$ at the probing stage. $\Gamma_{re}$ is the measured spin relaxation rate of the FID signal and $\phi$ is the phase of the FID signal. By fitting the FID signal with Eq. (4), we can extract the initial amplitude of the FID signal, the spin relaxation rate, and the Larmor frequency.
In order to estimate the sensitivity of our magnetometer, we apply a frequency counter (53230A, Keysight) which measures the frequency of the signal based on the zero-crossing detection method. The reciprocal of time between two zero-crossing points becomes the Larmor frequency.
Figure 4: (a) FID amplitude as a function of LD current DC offsets and modulation depths. The results are separated into four regions divied by two dashed line (the boundaries where modulation includes mode-hop current) and straigt line(the mode-hop current at 67.1 mA). (b) Semi-theoretical result. $S_{0}$ is calculated based on the optical pumping profile shown in the top pannel. The optical pumping rate profile is extracted by subtracting minimum relaxation rate from the measured relaxation rate shown at Fig. 3 (c). (c) Semi-theoretical result based on the step-function like optical pumping profile. (d) FID signals at four regions. The insets show 1-period temporal pumping rate profiles extracted from the top pannel of (b). The modulation depth is fixed to 1.8 mApp.
IV Experimental results and discussion
Changes in the LD injection current lead to concurrent changes in laser beam intensity and detuning, consequently affecting both the Faraday rotation angle and the spin relaxaton rate. In order to examine these parameters as a function of the LD injection currents, we obtained FID signals by chainging the LD current DC offset at the probe stage while current modulation at the pumping stage is fixed. Because the spin polarization remains unchanged (due to fixed current modulation at the pumping stage), the FID signal reflects the atomic response on the LD current DC offset. Fig. 3 (a) depicts the absorption spectra of 87Rb in the atomic vapor cell. Fig. 3 (b) shows the laser frequency detuning as a function of LD injection currents. A huge shift on the laser frequency occurs around 67.1 mA. The relative size of Faraday rotation angles and transverse spin realxation rates as a function of LD injection currents is shown at Fig. 3 (c). At the current lower than the mode-hop current, the laser becomes near-resonant so that a large Faraday rotation angle and a fast spin relaxation are induced. The LD current higher than the mode-hop current, conversely, shifts the laser frequency into off-resonance, leading to a relatively small Faraday rotation angle and a smaller spin relaxation rate. The maximum difference in spin relaxation rates between on- and off-resonance was approximately 1036 /s, which represents the maximum optical pumping rate. The minimum spin relaxation rate in off-resonance is 2392  /s when the $1/e^{2}$ beam diameter is 10 mm which is the internal size of the atomic vapor cell. When a beam diameter is small, Rb-Rb spin exchange collisions act as the main spin relaxation mechanism. A polarized atomic spin in the boundary between the bright region and the dark region is easily depolarized due to a spin exchange collision with an unpolarized atomic spin from the dark region. With a beam diameter of 1.5 mm, the minimum spin relaxation rate was measured as 4429  /s which is very close to 4304  /s, the 87Rb spin-exchange relaxation rate at the cell temperature of 98 C gibbs1967spin . It means that the optical pumping can be effectively turned off by using laser mode hopping. Fig. 3 (c) shows the optical power of the laser beam after passing through the atomic vapor cell. The absorption is minimized at the off-resonance condition (laser current above the mode-hop current), leading to a higher FID amplitude $V_{0}$ in spite of the slight degradation of the Faraday rotation angle.
We measured the initial amplitude of the FID as a function of the LD current DC offset and the current modulation depth as shown in Figure. 4. Fig. 4 (a) shows the measured relative FID amplitude. We divide Fig. 4 (a) into the four regions depending on whether the current modulation in the pumping stage includes the laser mode-hop current and whether the current in the probing stage is larger(or smaller) than the laser mode-hop current. The black dashed lines in Fig. 4 (a)-(c) indicate the boundary where the current modulation in the pumping stage contains the laser mode-hop current. The black solid line depicts the mode-hop current, 67.1 mA. The laser frequeny in the probing stage is off-resonant (or on-resonant) in the region to the right (or left) of the black solid line. We note that the experimental data shown in Fig. 4 (a) reflects the combined effect of the paramagnetic Faraday rotation, beam absorption and the discontinuous optical pumping.
In the region 1, current modulation in the pumping stage excludes the laser mode-hop current so that the continuous optical pumping profile with a small $R_{1}$ causes the small $S_{0}$. Additionally, the current in the probing stage brings the laser frequency into on-resonance, resulting in a small beam intensity due the strong absorption. These lead to the small FID amplide as depicted by the black line in Fig. 4 (d). In the region 2, current modulation in the pumping stage does not contain the mode-hop current like the region 1. However, $R_{1}$ of the region 2 is smaller than region 1 as shown in the insets of Fig. 4 (d). Hence, $S_{0}$ of the region 2 is expected to be smaller than that of the region 1. On the other hand, the laser frequency at the probing stage is off-resonant so that the spin relaxation rate is much smaller than that of region 1. The resulting FID signal is depicted by the green line in Fig. 4 (d), whose amplitude is slightly smaller than region 1 (black line). In the region 3, the mode-hop current is included in the current modulation, resulting in a highly discontinuous modulation profile. A large $S_{0}$ is produced and a large Faraday rotation occurs at the probing stage. In this region laser frequency at the probing stage is on-resonant. The absorption of the laser beam, therefore, results in a degradation of FID amplitude, as shown in red line in Fig. 4 (d). The maximum FID amplitude is shown in the region 4, as depicted by the blue line in Fig. 4 (d). A large $S_{0}$ is induced because the current modulation including the laser mode-hop current realize a discontinous optical pumping profile with a large $R_{1}$. The off-resonant beam minimizes absorption at the probing stage so that a large FID amplitude is induced compared to region 3. Moreover, the spin relaxation rate of the region 4 is slower compared to that of region 3, which is beneficial in terms of magnetometer sensitivities.
To compare the experimental data with the theory, Eq. (2) is calculated using $R_{0}$ and $R_{1}$ extracted from the $R(t)$ profile shown at the top panel of Fig. 4 (b) and (c). The $R(t)$ profile in the top panel of Fig. 4 (b) is obtained by subtracting the minimum relaxation rate from the measured spin relaxation rate shown at Fig. 3 (c). Fig. 4 (b) shows the calculated relative FID amplitude using the triangular function like $R(t)$ profile in the top panel of Fig. 4 (b). Fig. 4 (c) is derived from the fictious step-function like $R(t)$ profile given in the top panel of Fig. 4 (c). Both the experimental result and the semi-theoretical results seem to have similar tendency. But the details are slightly different from each other. Figure 5 shows maximum FID amplitudes as a function of modulation depth and the LD current DC offsets with the maximum FID amplitude in the experiment (Fig. 4 (a)) and two semi-theoretical calculations (Fig. 4 (b) and (c) ). Because two semi-theoretical results and the experimental result use the same Faraday rotation profile and the same laser beam absorption profile, the differences between them are mainly attributed by discontinuous optical pumping.
Figure 5: (a) Maximum FID amplitude as a function of modulation depth. The black dots represent the experimental data, derived from Fig. 4 (a). The red line and the blue line depict the semi-theoretical results extracted from Fig. 4 (b) and Fig. 4 (c), respectively. The experimental data(black dots) is more robust to the change of modulation depth compared to the semi-theoretical result(red line) where the optical pumping profile is a triangular function. The semi-theoretical result (blue line) where the optical pumping profile is a step function shows that the maximum ampltiude of FID increases with increasing modulation depth. (b) LD current DC offset with the maximum FID amplitude as a function of modulation depth.
Comparing experimental data (black dots) and semi-theoretical result with the measured triangular function like optical pumping profile (red line) in Fig. 5 (a), the maximum FID amplitude decreases moderately as the modulation depth grows while $S_{0}$ derived by Eq. (2) fastly decreases. It implies that the measured triangular function like optical pumping profile used in the top panel of Fig. 4 (b) does not fully describe the mode-hopping asssited optical pumping. This discrepancy between the semi-theoretical result and the experimental result needs to be studied in the future. Here, we describe its potential origins. First, we note that the optical pumping rate profile shown in Fig. 4 (b) is calculated from the transverse spin relaxation rates extracted from the FID signals as a function of the LD current DC offset in the probing stage. The FID signals are affected by the constant laser frequency. However, because the laser frequency sweeps in the pumping stage, the optical pumping rate in the pumping stage is slightly different from the optical pumping rate derived from the FID signals as a function of the LD current DC offset in the probing stage. In particular, when the ground hyperfine structure is considered, the optical pumping rates of the $F_{g}=1$ state and the $F_{g}=2$ change with times because of the laser frequency sweeping. The transverse spin along the $\hat{x}$-direction can be represented by $\langle S_{x}\rangle=(\langle F_{x}^{I+1/2}\rangle-\langle F_{x}^{I-1/2}\rangle)/(2I+1)$ where $I$ represents the nulcear spin, $I=3/2$ for 87Rb. To maximize spin polarization $S_{0}$ of 87Rb, the $F_{g}=2,~{}m_{F_{g}}=2$ state (or $m_{F_{g}}=-2$) should be mainly populated. For this, the hyperfine repumping of the $F_{g}=1$ state is required li2022repumping ; schultze2015improving . Due to the large pressure broadening of a order of GHz, the laser at the pumping stage pumps both the $F_{g}=1$ and $F_{g}=2$ states with different optical pumping rates. The different time-varying optical pumping rates for the $F_{g}=1$ and the $F_{g}=2$ states have to be taken into account to describe the mode-hopping assisted optical pumping more accurately. Second, the spatial inhomogeneity of the optical pumping profile needs to be considered. Due to light absorption, the optical pumping beam attenuates along the beam propagation direction so that the optical pumping profile is spatially inhomogeneous. The measured optical pumping profile shown in the top panel of Fig. 4 (b) is a kind of spatially averaged values. The analysis considering the spatially inhomogeneous optical pumping profile will give more accurate results. Additionally, we obtain the semi-theoretical result with the step function like optical pumping profile as depicted by the blue solid line in Fig. 5 (a). In contrast to the experimental result and the semi-theoretical result with the triangular function like optical pumping profile, the relative FID amplitude increases with increasing modulation depth. Interestingly, the experimental data is between two semi-theoretical results. We can conjecture that the real optical pumping profile is between the triangular function and the step function.
We estimate the sensitvity of our magnetometer based on the laser mode-hopping assisted optical pumping. We measured Larmor frequencies at a repetiton rate of 1 kHz by using the commercial frequency counter. The measured Larmor frequencies were converted into the magnitude of the magnetic field by diving the Larmor frequencies by the gyromagnetic ratio of 87Rb. Finally, the amplitude spectral density (ASD) of the magnitude of the magnetic field was obtained. The inset of Figure 6 (a) shows averaged ASD of 1000 shots of the FID signal when the beam intensity after passing through the atomic vapor cell was 0.9 mW. Fig. 6 (a) shows the sensitivity of our magnetometer obtained by measuring almost 5 sec. We acheived a sensitivity of 5.57 pT/Hz1/2, mainly limited by the performance of the frequency counter. We note that a sensitivity of 3.77 pT/Hz1/2 was obtained in the laser power of 1.4 mW where the FID signal was partially a square-like wave due to the saturation of the BPD.
The measured sensitivity is mainly limited by the amplitude noise in the frequency counter. Considering a damped oscillation with a RMS amplitude noise $\delta y$, the amplitude noise is converted into a frequency error via the slew rate at the threshold. When the frequency is estimated by averaging the time intervals between the zero-crossing time stamps (mean frequency model of frequency counter), the deviation of the magnetic field error $\rho_{fc}$ averaged over the measurement time of $T_{m}$ is written by
$\rho_{fc}=\frac{\delta y}{\gamma V_{0}\sqrt{f_{BW}}T_{m}}\sqrt{1+e^{2\Gamma_{re}T_{m}}}.$ (5)
Here, $V_{0}$ is the FID ampliutde and $f_{BW}$ is the bandwidth of the magnetometer. The total rms amplitude noise $\delta y$ is given by $\delta y=\sqrt{\delta^{2}y_{fc}+\delta^{2}y_{ph}}$ where $\delta y_{fc}$ is the amplitude noise of the frequency counter, 3500 $\mu$$\textrm{V}_{\textrm{rms}}$ and $\delta y_{ph}$ is the RMS photon shot-noise. The calculated magnetic field noise is shown at Fig. 6 (b) as the black dashed line which shows good agreement to the measured magnetometer noise floor.
Our frequency measurement using the commercial frequency counter is not optimal. The potentially achievable sensitivity can be estimated by the Cramer-Rao lower bound (CRLB) calculated by a shot of FID signal gemmel2010ultra . The Cramer-Rao lower bound (CRLB) is the sensitivity reference for an atomic magnetometer based on measurement of the Larmor frequency of the decaying FID signal. The CRLB is given by
$\rho_{CRLB}=\frac{\sqrt{12C_{w}}\rho_{\omega}}{\gamma V_{0}\sqrt{f_{BW}}T_{m}^{3/2}}.$ (6)
where $\rho_{\omega}$ is the ASD level of the BPD signal at the Larmor freuqency. The coefficient $C_{\omega}$ which depends on the transveral spin relaxation rate $\Gamma_{re}$ is given by
Figure 6: (a) Noise spectral density of our magnetometer. The black line is the magnetometer noise floor of 5.57 pT/Hz1/2 which is averaged over 10 Hz to 500 Hz when the commercial frequency counter is applied. The red dashed line shows a CRLB of 191 fT/Hz1/2 which is calculated from Eq. (6). The inset shows the ASD of the FID signal which is used for the calculation of the CRLB. The red line depicts the ASD when $\Delta w=|w-w_{0}|\simeq 0$. The black line depicts the ASD when $\Delta w=100$ kHz, which determines $\rho_{\omega}$. A small 85Rb peak is shown at 143 kHz. The resolution bandwidth of the ASD is 3.9 kHz which is limited by the magnetometer bandwidth. The low frequency ASD includes an internal electronic noise arising from the spectrum analyzer(EXA Signal Analyzer N9010A). The photon shot-noise level is depicted by the dashed line for comparison. The beam power after passing through the atomic vapor cell was 0.9 mW. (b) Magnetometer noise floor and CRLB as a function of beam power after passing thorugh the atomic vapor cell. The black dots depict the measured magnetometer noise floor using the commerical frequency counter. The black dashed line shows the calculated noise level from mean frequency model (Eq. (5)). The red dots show the CRLB calculation based on the ASD measurement, which represent the reachable sensitivity of the magnetometer (Eq. (6)). The red dashed line shows the photon shot-noise limited CRLB calculation.
$C_{w}=\frac{e^{2/r}-1}{3r^{3}\textrm{cosh}(2/r)-3r(r^{2}+2)}.$ (7)
where $r=1/(\Gamma_{re}T_{m})$. The CRLB calculated by Eq. (6) represents the non magnetic-oriented noise induced sensitivity. When the beam intensity after absorption is 0.9 mW, $\Gamma_{re}=2358$ /s, $T_{m}=460$ $\mu$s, $\rho_{\omega}=2.6$ $\mu$V/Hz1/2, $V_{0}=8.5$ V, and $f_{BW}=500$ Hz, we obtain the CRLB of 191 fT/Hz1/2. Fig. 6 (b) shows the measured magnetometer noise level and CRLB for various laser power after absorption throughout the vapor cell. The red points show the CRLB calculation assuming that $\rho_{\omega}$ is limited to the photon shot-noise which is given by $\sqrt{eG^{2}RI_{0}}$. Here, $G$ is the transimpedence gain, $R$ is the reponsivity of the BPD and $e$ is the classical electron radius. The CRLB calculation shows that our magnetometer is potentially capable of reaching a sensitivity of a sub pT/Hz1/2 by revising the frequency counting method.
We note that the frequency shift of 20 GHz and the laser frequency in the pumping stage which is moderately detuned from the resonance are not optimal. In order to mainly populate the $F_{g}=2,~{}m_{F_{g}}=2$ state for maximizing $S_{0}$, the on-resonance laser frequency in the pumping stage should be close to the $F_{g}=1,2$ resonances. Also the laser freuqency needs to be far-detuned from the resonance periodically for synchronous optical pumping to obtain a large optical pumping modulation depth. A careful laser cavity adjustment in laser diode fabrication, leading to an optimal frequency shift of mode-hopping and the on-resonance laser frequency very close to the resonance, will be helpful to increase the sensitivity of magnetometer. In addition, multi-pass probing will increase the sensitivity because $V_{0}$ is proportional to the interaction length li2011optical .
V Conclusion
We have demonstrated a single-beam all-optical pulsed atomic magnetometer enhanced by laser mode-hopping. Optical pumping assisted by laser mode-hopping, which leads to a discontinuous pumping profile, effectively enhances the transverse spin polarization. Furthermore, mode-hopping offers a convenient method for setting probing stage beam off-resonant which suppress the absorption and maintain large FID signal. We have reached a sensitivity of 3.77 pT/Hz1/2 which is sufficent for applications of magnetic anomaly detection in earth’s magnetic field. The sensitivity of our magnetometer seems to be limited by the performance of the frequency counter. In order to estimate the potentially achievable sensitivity, CRLB due to the non-magnetic oriented noise was calculated, which was 191 $\textrm{fT}/\sqrt{\textrm{Hz}}$ in our experimental condition. Considering the calculated CRLB, we expect that our magnetometer can reach a sensitivity of a sub pT/Hz1/2 which is similar to the sensitivity limited by photon shot-noise. A laser cavity optimization considering laser mode-hopping in fabrication of a DBR LD may further improve the sensitivity by ehancing optical pumping. Our approach based on laser mode-hopping can be applied to miniaturization of all-optical atomic magnetometers with sub-pT/Hz1/2 sensitivities.
VI Acknowledgments
This work was supported by the Agency for Defense Development Grant funded by the Korean government (No. 915098102).
VII Appendix
VII.1 Consideration on Faraday Rotation of Elliptically Polarized Beam
The spin polarization formed assisted by mode-hopping enters the probing stage where elliptically polarized light undergoes optical Faraday rotation, then measured by a polarimetry. In a linear polarization basis ($\hat{e}_{H},\hat{e}_{V}$), the elliptical polarization state is represented as $(\cos(\theta),\sin(\theta)\exp(i\phi))$. Here, $\theta$ is the tilt angle, and $\phi$ is the retardation angle. The change in polarization state of the beam after passing through the atomic vapor can be expressed by the matrix $M_{c}=\begin{pmatrix}{1+e^{i\Phi}}&{i(1-e^{i\Phi})}\\ {i(e^{i\Phi}-1)}&{1+e^{i\Phi}}\end{pmatrix}$
where $\Phi=|s|\Phi_{0}(t)$ is the Faraday rotation angle. Here, $\Phi_{0}(t)$ is damped oscillation in our experiment and $|s|$ is the size of photon spin vector which is calculated by $\sin(2\theta)\sin(\phi)$. The matrix representation of the HWP is $M_{h}=\begin{pmatrix}{\cos(2\theta_{h})}&{\sin(2\theta_{h})}\\ {\sin(2\theta_{h})}&{-\cos(2\theta_{h})}\end{pmatrix}$, so the final polarization state of the elliptically polarized beam measured by the polarimeter is $M_{h}\cdot M_{c}\cdot(\cos(\theta),\sin(\theta)\exp(i\phi))$. The balanced signal is then calculated by the difference between squares of each orthogonal linear polarization components of the beam: $4I_{0}[\cos(4\theta_{h}+\Phi)\cos(2\theta)+\cos(\phi)\sin(4\theta_{h}+\Phi)\sin(2\theta)]$. Here, $I_{0}$ is the intensity of the beam after absorption throughout the vapor cell. When $1\gg\Phi$, the oscillating signal simplifies to
$I_{B}\approx I_{0}[\cos(4\theta_{h})\cos(2\theta)+\sin(4\theta_{h})\sin(2\theta)\cos\phi\\ +|s|\Phi_{0}(t)(-\sin(4\theta_{h})\cos(2\theta)+\cos(4\theta_{h})\sin(2\theta)\cos\phi)].$ (8)
Figure 7: The polarization state of incident beam(upper plot) and relative FID amplitude(bottom plot) as a function of QWP angle. When the QWP angle is 25 where relative FID amplitude is maximized, $\theta$ is 30 and $\phi$ is 129 . The red line shows $\sin(4\theta_{q})$ behavior to the QWP angle $\theta_{q}$. Blue line illustrates Eq. (8) calculated based on the polarization state on the top plot.
By adjusting the HWP angle $\theta_{h}$, the baseline component of the balanced signal can be eliminated. The amplitude coefficient proportional to $\Phi_{0}(t)$, simultaneously, is maximized.
Previous study has shown that signal amplitude in single-beam SERF magnetometers follows a $\sin(4\theta_{q})$-dependence on QWP angle $\theta_{q}$shah2009spin . Our all-optical single beam pulsed magnetometer shows similar behavior under the bias field of 14.3 $\mu T$ which is comparable to the Earth’s ambient magnetic field. The top plot of Figure. 7 shows the polarization state of the beam before entering the vapor cell as measured by a polarimeter(Thorlabs PAX 5710IR1) at various QWP angles. The bottom plot shows the relative FID amplitude measured at different QWP angles. The red line illustrates $|\sin(4\theta_{q})|$ behavior, while the blue line represents equation (8), which is calculated based on the polarization state shown in the upper plot. HWP angle $\theta_{h}$ is adjusted to compensate signal baseline to be zero for each measurement.
We determined the optimal QWP angle for single-beam all-optical magnetometer to be 25 with a 133 retardation, suggesting an elliptically polarized beam as the most effective polarization state. The elliptical polarization strikes a balance between the most efficient pumping of circular polarization and the most effective Faraday rotation that occurs with linear polarization. Note that for FID measurements, compensating baseline through HWP angle adjustment is essential. Without baseline compensation, the FID amplitude may decrease, and if the beam intensity is too high, the baseline component may saturate the entire signal. To avoid signal saturation on the BPD signal, in our measurement, the beam intensity in front of the vapor cell is kept at 1.43 mW during this measurement.